OpenStudy (anonymous):

what would this be in a fraction in simplest form::: 2nickles 1 dime to 10 pennies

Parth (parthkohli):

10 penny ===> 10 cents 2 nickels and 1 dime ===> 10 + 10 = 20 cents

Parth (parthkohli):

Ratio is the same as fraction. So \(10:20 = {10 \over 20}\)

OpenStudy (anonymous):

(2/20) + (1/10 ) + (10/100) (10/100) + (10/100) + (10/100) = 30/100 = 3/10

OpenStudy (anonymous):

well 2 nickles and 1 dime is 20 cents and 10 pennies is 10 cents it says 2 nickles 1 dime to 10 pennies so wouldnt it be (in ratio form) 20:10 and in fraction: 20/10 and then 2/1 in simplest form?

OpenStudy (anonymous):

yes 2:1

OpenStudy (anonymous):

so in fraction, would it be 2 over 1 (2/1) an improper fraction

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks!